OpenStudy (angel123):

PLZ HELP!!! will give medal and become fan!! Rumor has it that a woman can drink blueberry tea to increase the likelihood of having a girl baby to three times that of having a boy. a) If a woman drinks the tea daily what are the chances that, out of five children, there are at least 2 girls? b) What is the expected number of girls in 5 children?

zepdrix (zepdrix):

@TheSmartOne @mathstudent55 @agent0smith I dunno, sorry :c

OpenStudy (mrnood):

'rumour has it' is not a statistical statement. The chances of having a girl are approximately the same as having a boy, regardless of the rumour

OpenStudy (angel123):

yea but how would u solve it?

OpenStudy (mrnood):

if you assume the probability of girl = possibility of boy then this is th esame as tossing a coin you are studying combinations and permutations look up the formula for tossing at least 2 heads from 5 throws. That is the correct answer (HOWEVER my guess is that the question expects you to accept the ridiculous rumour...)

OpenStudy (angel123):

we did learn combination and permutation but we are learning probability distributions right now so..

OpenStudy (mrnood):

they are linked this is binomial probability

OpenStudy (agent0smith):

"chance of a girl baby to three times that of having a boy" so chance of a girl must be 0.75, chance of a boy 0.25 a) If a woman drinks the tea daily what are the chances that, out of five children, there are at least 2 girls? It's binomial probability. Hopefully you're familiar with it. \[\Large P(X = r) = \left(\begin{matrix}n \\ r\end{matrix}\right) p^r q^{n-r}\]p=0.75 q=0.25 n=5 You need to find \(\large P(X \ge 2)\) You can use some things to simplify the work\[\large P(X \ge 2) = 1 - P(X \le 1)\]where\[\large P(X \le 1) = P(X=0) + P(X=1)\]

OpenStudy (mrnood):

I stress again 'rumour has it' does NOT state the Pg=3Pb

OpenStudy (agent0smith):

Lots of questions ask us to accept some nonsense and use it anyway.

OpenStudy (angel123):

so would r=3 then solve for p(x=3)

OpenStudy (agent0smith):

Depends what you're doing. If you're doing \[\large P(X \ge 2) = P(X=2) +P(X=3) +P(X=4) +P(X=5)\]then yes. But hopefully you see why that is more work than finding \[\large P(X \ge 2) = 1 - P(X < 2) \]or\[\large P(X \ge 2) = 1 - [P(X=1)+P(X=0) ]\]

OpenStudy (angel123):

oh okay makes sense thanks alot !!! u helped alot

OpenStudy (angel123):

how about b)

OpenStudy (agent0smith):

Really easy. b) What is the expected number of girls in 5 children? Multiply probability of a girl by the total children.

OpenStudy (angel123):

nvm i got it

OpenStudy (angel123):

thank you